So i need to find the surface area of a half of a cylinder. I've tried everything i know how to do and even looked online and everything, but i cannot find the surface area without a width, and on the problem there is no width.
the length is 5
the diameter is 4

Answer by oberobic(2304)

The surface area of 1/2 a cylinder is 1/2* surface area of the whole cylinder.
.
L = the length of the cylinder = 5
.
D = diameter = 4
.
Surface area = circumference * L
.
Surface area = diameter * pi * L
.
Surface area = 4 * pi * 5 = 20 pi
.
1/2 surface area = 1/2 * 20 pi = 10 pi
.
Done.